015 10.0 points The Pampas lowlands of Argentina extending into Uruguay and adjacent areas of Paraguay are notorious for the number of tornadoes that occur, on average 3 per week! In a given 4 week period, what is the probability that more than 4 occur?
Answer #1

This follows a Poisson Distribution,

Average,  λ = 3 per week

In a 4 week period, Probability that more than 4 occur is equivalent to-

If X be the number of tornadoes per week

P(X>1) is what is required

1- P(X=0)-P(X=1)

In Poisson distribution, P(X=x) = (e^- λ * λ^x)/x!

At x=0, P(X=0) = {e^-3 * 3^0}/0! = 0.049

At x=1, P(X=1) = {e^-3 * 3^1}/1! = 0.149

Now, P(X>1) = 1-0.049-0.149 = 0.8016

Hope this helps and Please don't forget to like the answer :)